OpenStudy (anonymous):

what is the acceleration of a body moving with constant acceleration covers the distance between two points 60m apart in 60sec? its velocity as it passes the second point is 15m/sec

OpenStudy (kira_yamato):

Constant acceleration implies that v(t) is a linear function am I right?

OpenStudy (anonymous):

yes

OpenStudy (kira_yamato):

So what does this tell you about your s(t)?

OpenStudy (kira_yamato):

Hint: Use integration.

OpenStudy (anonymous):

a = dv/at??

OpenStudy (kira_yamato):

If a = v'(t), can we say that v(t) = u + at for some constant u?

OpenStudy (kira_yamato):

Recall: Integration is the reverse of differentiation.

OpenStudy (anonymous):

a = d^2s/dt^2??

OpenStudy (kira_yamato):

Yep. But let's see it in the integral form instead. ^^ So\[v(t) = \int\limits a dt\] and \[s(t) = \int\limits v(t) dt\]

OpenStudy (kira_yamato):

So you know a is a constant, so v(t) = u + at, where u is the arbitrary constant of integration right?

OpenStudy (kira_yamato):

The question states that "its (the body's) velocity as it passes the second point (at t = 60) is 15m/s. So we know that v(60) = 15. Solve for u in terms of a.

OpenStudy (anonymous):

oh.. ok ok sir im still kinda confuse but ill try to solve im not really good at integral calculus sorry

OpenStudy (kira_yamato):

It's alright ^_^ v(t) = u + at v(60) = u + a(60) 15 = u + 60a u = 15 - 60a v(t) = 15 - 60a + at

OpenStudy (anonymous):

sir the V is = to 15?

OpenStudy (kira_yamato):

\[s(t) = \int\limits{(15 - 60a + at)}dt = 15t + 60at + 0.5at^2 + c\] where c = arbitrary constant of integration. Let's use s(0) = 0 15(0) + 60(0)a + 0.5(0^2)a + c = 0 c = 0 Hence \[s(t) = 15t + 60at+0.5at^2\] Also, s(60) = 60 15(60) + 60(60)a + 0.5(60^2)a = 60 390a = -840 a = -84/39

OpenStudy (kira_yamato):

Note that v = 15 only at the point where t = 60.

OpenStudy (anonymous):

oh.. ok so the first v you use is at rest right?

OpenStudy (anonymous):

uhh.. sir according here the answer for a = 1.67

OpenStudy (anonymous):

@Kira_Yamato sir?

OpenStudy (kira_yamato):

I don't think so... Because you don't have the value of v(0).

OpenStudy (shamil98):

v (final velocity) =15m^s s (displacement) =60m t (time taken) = 60s acceleration =a, initial velocity =u, v =u+ at 15m^s = U + a(60s) a = (15m^s -u)/60s v^2 - u^2 =2as (15m^s)^2 - u^2 =2*60m*(15 m^s - u)/60s) this is what im at so far...

OpenStudy (charlotte123):

there must be choices for dis question O_O

OpenStudy (shamil98):

its a quadratic equation, just solve for then plug in the value of u into the previous equation and solve for a.

OpenStudy (kira_yamato):

So it had nothing to do with integral calculus?

OpenStudy (shamil98):

nope..

OpenStudy (shamil98):

this is physics lol

OpenStudy (charlotte123):

wait, is it 60 seconds or 6? ^_^ @Shin_

OpenStudy (charlotte123):

oh dang hes offline .-.

OpenStudy (charlotte123):

http://answers.yahoo.com/question/index?qid=20070908094909AAOBpps <--- similar @_@

OpenStudy (charlotte123):

http://answers.yahoo.com/question/index?qid=20090716031821AADpzfB <--- and this one too :3

OpenStudy (charlotte123):

I am such a coconut .-.

OpenStudy (kira_yamato):

I have all along been using calculus to solve for kinematics that I forgot my equations of motion... :(

OpenStudy (charlotte123):

@Kira_Yamato I remember doing that years ago XD It was for science class @_@ unless this is a different subject, den I would be X_x

OpenStudy (shamil98):

i have no idea how one would solve for this with calc lol, so maybe o.o, but yeah remember your equations for motion :P

OpenStudy (charlotte123):

@shamil98 that is what I was thinking XD

OpenStudy (charlotte123):

oh and this one :D

OpenStudy (kira_yamato):

(15)^2 - u^2 =2*60*(15 - u)/60) 225 - u^2 = 30 - 2u u^2 - 2u - 195 = 0 u = 15 or u = -13 Sub. u = 15 into a, a = (15 - 15)/60 = 0 Sub. u = -13 into a, a = (15 + 13)/60 = 7/15

OpenStudy (charlotte123):

@Kira_Yamato YES :D I suck XD

OpenStudy (kira_yamato):

How I used it was this: 

OpenStudy (charlotte123):

I don't know what that is, but great job @_@ XD

OpenStudy (shamil98):

v^2 - u^2 =2as plug them in now lol let the velocity(u) at the start be 0.

OpenStudy (charlotte123):

I used the internet *facepalm* >.>

OpenStudy (shamil98):

Same here char .-. , my physics knowledge is very limited..

OpenStudy (charlotte123):

@shamil98 something we are the same on XD